Counting Permutations that Avoid Many Patterns

Yonah BIERS-ARIEL, Haripriya CHAKRABORTY, John CHIARELLI, Brian EK,

Andrew LOHR, Jinyoung PARK, Justin SEMONSEN, Richard VOEPEL, Mingjia YANG, Anthony ZALESKI, and Doron ZEILBERGER

Abstract:

0. Maple Package and Sample Output

This article is accompanied by a Maple package, C27.txt, that is available, along with six input and output files, from

http://www.math.rutgers.edu/~zeilberg/mamarim/mamarimhtml/expmath27.html .

8. References

Doron Zeilberger, Department of Mathematics, Rutgers University (New Brunswick), Hill Center-Busch Campus, 110 Frelinghuysen Rd., Piscataway, NJ 08854-8019, USA;

DoronZeil at gmail dot com ; http://www.math.rutgers.edu/~zeilberg/ .

To be published in The Personal Journal of Shalosh B. Ekhad and Doron Zeilberger

(http://www.math.rutgers.edu/~zeilberg/pj.html), N. J. A. Sloane’s Home Page
(http://neilsloane.com/), and the arxiv.org .

Dec. 24, 2016